hi guys, we are really not getting how a algebraic equasion from a word
problem, eg
A woman is 6 yrs older than three times her daughters age. Their ages total
46. What are their ages?
Let x = daughter's age
Let 3x = mother's age
x + 3x=46 11.5 +3(11.5)=?46
4x =46
x 11.5 46=46
3x = 3(11.5)
3x= 34.5 woman's age 11.5 = daughter's age
Are we right or have we just not got it need some hints

Answer by psbhowmick(878)

Let daughter's age = x yrs.
Then, mother's age = (3x + 6) yrs.
x + (3x + 6) = 46
4x = 46 - 6 = 40
x = 40/4 =10

3x+6 = 36 = mother's age; x = 10 = daughter's age
This is the correct answer!
Mother's age is (3x + 6) yrs. NOT 3x yrs.